The impedance of a circuit consist 0f 3 ohm resistance and 4 ohm reactance. The power factor of the circuit is :- a) 0.4 b) 0.6 c) 0.8 d) 1.0

Grade:

5 Answers

suvidhi mehta
6 years ago
answer b i.e. 0.6
hope its correct.
suvidhi mehta
6 years ago
then what is the answer my friend..
 
i think i i help you if you provide the correct  answer....
 so sorry for inconvenience
Sarthak Khandelwal
6 years ago
What`s the solution answer is correct.
suvidhi mehta
6 years ago
power factor for an AC ckt is given by R/{R + (XL – Xc)}^1/2
where R is resistance 
XL is inductive reactance
Xc is capacitive reactance
Hope it helps you.
Shajid
5 years ago
Power factor=R/Z
Z=(R^2+XL^2)^1/2=(9+16)^1/2=5
Power factor=3/5=0.6 
where R is resistance 
XL is inductive reactance
Xc is capacitive reactance
Hope it helps you.